OpenStudy (anonymous):

how determine the value of x and y such that points (1,2,3),(7,3,1),(x,y,2) are collinear

OpenStudy (amistre64):

the ratio of parts are proportional

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

wat do u meam parts are proportional

OpenStudy (anonymous):

mean

OpenStudy (amistre64):

bad idea, i was thinking of parralel vectors: create a line from the known parts r = (1,2,3) + t(7,3,1) x = 1 + 7t y = 2 + 3t z = 3 + t determine (x,y,z) = (x,y,2)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

2 = 3+t ; t = -1 solve for x and y with t=-1

OpenStudy (anonymous):

x=-6 t=-1 y=-1

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

hi my answers are wrong

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

se ΔPQR to answer the question. What is tan P?

OpenStudy (phi):

what you want for the equation of the line is a vector that points in the direction of the line one way to find that vector is subtract 2 points on the line: V= (7,3,1)- (1,2,3)= (6,1,-2) the equation of the line is some point on the line, plus some amount of distance in the direction of V for example: P= (1,2,3) + t ( 6,1,-2) if we set t= 0 we get point (1,2,3) if we set t=1, we get point (7,3,1) to find the x,y values to so that point (x,y,2) is on the line: (x,y,2)= (1,2,3) * t (6,1,-2) we need to find the correct value for "t" one way: set the 3rd components equal: 2 = 3+ t*-2 we find t= 0.5 now you can find x and y: x= 1+ 0.5*6= 4 y= 2+0.5*1= 2.5 test: (1,2,3)+ 0.5*(6,1,-2)= (1,2,3)+ (3, 0.5, -1)= ( 4, 2.5, 2)

OpenStudy (anonymous):

ok 

OpenStudy (anonymous):

thanks